### v1

@@ II.2 NHQM methods
Figure 1: A resolvent matrix element $\mathcal{A}(E)$ in the complex $E$ plane with fixed sources. Here, we explore the off-shell behavior of the non-Born term of a $T$-matrix. The two rows show the real and imaginary parts of the amplitude. From left to right, three different calculations are shown. The first two are based on bound-state-like calculations, where we enforce a Dirichlet boundary condition on the relative motion, forcing the wave functions to be zero at $r=R_{IR}$; the third is the exact calculation, corresponding to the $R_{IR}\to\infty$ limit.
To illustrate the NHQM methods, including the one developed here, we use a two-body toy model with finite-range interactions. The toy system mimics a two-nucleon system in $s$-wave (see Sec. IV for details), for which we compute $\mathcal{A}$ in various ways. The sources are chosen such that $\mathcal{A}$ corresponds to the non-Born term in the scattering $T$-matrix, previously discussed as one option in Sec. II.1. Here, we focus on the generic features of the numerical results without dwelling on the details.
@@
The figure shows the real (top row) and imaginary (bottom row) parts of various $\mathcal{A}(E)$s in the complex $E$ plane; from left to right, the first two columns are from the $\mathcal{A}$ calculations in which we plug finite $H$ matrices into Eq. (2). To get such $H$ matrices, we enforce a boundary condition such that all the states are zero at $r=R_{IR}$ (10 and 50 fm in the 1st and 2nd column), a particular long-distance (infrared, IR) regulator, while a series of short-distance (ultra-violet, UV) regulators [^5] For example, in the momentum space calculations, the regulator could be a cut-off on the maximum momentum. are tested such that the results shown in the figure converge. The last column shows the exact results, essentially $\mathcal{A}$’s limit at $R_{IR}\to\infty$.
The sharp features around the real axis, as seen in the first two columns, can be understood via the spectrum of the $H$ matrices. Specifically, the approximated $\mathcal{A}$ with finite $R_{IR}$ and UV cut-off is a finite sum of poles if we consider it as a function of $E$:
@@
$|\Psi_{i}\rangle$ is the $i$th eigenstate of $\begin{bmatrix}H\end{bmatrix}$ and the associated eigenenergy $E_{i}$ determines a pole location. Note here and later, we label the finite-matrix-based approximations of a quantity, e.g., $\mathcal{A}$, by $\begin{bmatrix}\mathcal{A}\end{bmatrix}$. Since $\begin{bmatrix}H\end{bmatrix}$ is kept Hermitian while we construct a basis to project $H$, the poles are only on the real axis.
@@
We have three observations: (1) above and below the real axis, the exact $\mathcal{A}$ (in the 3rd column) is smooth almost every, except the $E=0$ point where the function is not as well behaved. When crossing the positive real axis, the amplitude experiences a discontinuity. This is a reflection of $E=0$ being $\mathcal{A}$’s branch point and the positive real axis as the BCT; (2) when taking $R_{IR}\to\infty$ limit, the poles to the right of the $E=0$ point become more and more densely distributed and eventually merge into a BCT along the real axis; and (3) to the left of $E=0$, there are no densely distributed poles except possible isolated bound-state poles (absent in this toy model). $E=0$ is such a nontrivial location because it is the model’s continuum threshold, below which the kinetic phase space for the scattering eigenstates is zero and above which the phase space starts increasing from zero [^6] This can also be visualized by plotting $-\mathrm{Im}\,\mathcal{A}(E)$ with $\mathrm{Im}\,E\to 0^{+}$. The function is zero below the threshold and grows starting at the threshold. For example, see the 3rd panel in Fig. 15.. The three observations combined explain the origin of the $\mathcal{A}$’s branch point, its connection with the threshold—a property of the $H$’s spectrum, and the choice of the BCT along the real axis due to $H$’s hermiticity.
@@
The above discussions can be generalized to more complex situations where multiple thresholds (or channels) exist. There are the same number of eigenstate groups as the number of thresholds [^7] The eigenstates of the full Hamiltonian provide a complete basis to expand the full space. See Chap. 7 in Ref. Newton (1982). At each threshold, a new group of eigenstates emerges as initiated by the corresponding incoming channel.. Each group starts to contribute to $\mathcal{A}$ above the corresponding thresholds in a simple additive way, i.e., the sum over the intermediate state in Eq. (6) can be broken into individual groups associated with different channels. At the infinite-dimension limit, those starting points become branch points, following our reasoning in the single-channel case. In addition, it will be interesting and also possible to generalize these conclusions to the cases with long-range repulsive Coulomb interactions (needed in nuclear physics), in which the branch points are also essential singularities due to the Coulomb-barrier-induced Gamow suppression factors.
@@
To understand the failure of the finite matrix calculations, we return to the first two columns in Fig. 1. The regions with the largest errors are around the real energy axis, where the poles are located888 It is worth pointing out that the discretized BCT poles can be partially understood through quantization conditions for the corresponding eigenvalues. At least for simple systems, the continuum physics can be extracted at the poles’ locations using Lüscher-type approaches Luscher (1991); Busch et al. (1998); Stetcu et al. (2010); Rotureau et al. (2012); Zhang (2020); Zhang et al. (2020); Kievsky et al. (2012); Furnstahl et al. (2014); Shirokov et al. (2018); Yu et al. (2024). These methods also extract the continuum physics by computing bound-state-like problems.. That is, unfortunately, where the continuum physics observables can be physically measured. At the same time, the calculations potentially already converge in the other regions. This is the essential feature of the rational approximation Trefethen (2023), to which Eq. (6) belongs, applied to a function with branch points: poles are assigned to mimic branch cut so that the approximation works in the region away from the BCTs but not around the BCTs. Therefore, the cause of the continuum-physics difficulty is the Hermiticity of the $\begin{bmatrix}H\end{bmatrix}$, although Hermitian Hamiltonians have been used extensively in computing discrete bound states.
@@
Existing NHQM methods Reinhardt (1982); Moiseyev (2011) solve this problem by giving up $\begin{bmatrix}H\end{bmatrix}$’s Hermiticity so that the BCTs and the problematic region of the rational approximation are moved away from the real axis. As a result, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ can converge with finite $R_{IR}$ around the real energy region. Moreover, if the BCT is moved further away enough that the resonance poles, which were on the second Riemann sheet [^9] A second Riemann sheet is sometimes called as an unphysical sheet, which could lead to the wrong impression that it is not important. On the contrary, the function’s behavior in that part of the complex plane is directly related to the function’s behaviors relevant to observations. when the BCTs are defined as along the real axis, are now in the region enclosed by the old and new BCTs, and become the poles on the 1st Riemann sheet. Therefore, the resonance poles can now be seen simultaneously in $\mathcal{A}$ with bound state poles. Eq. (2) then dictates that the resonances, as the bound state, are part of the $H$’s eigenstates. In addition, all the eigenstates, including the bound and resonance states and the so-called discretized continuum (DC) states that give rise to the discretized BCT poles in $\mathcal{A}$, are integrable. Therefore, the NHQM methods enable us to compute the continuum physics from the bound-state-like calculations.
@@
The other solution, which we follow here, is based on the observation that $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$ can converge to a certain required precision when the dimension of the $\begin{bmatrix}H\end{bmatrix}$ and the size of $|\mathrm{Im}\,(E)|$ are large enough. As shown in Fig. 1, with $R_{IR}=10$ fm, the computed $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ with $|\mathrm{Im}\,E|>10$ MeV already converges nicely to the exact results. This statement applies to general cases that the wave function solutions in Eqs. (3) and (4) for complex $E$ and compact sources satisfy the bound-state-like boundary conditions Efros (1985a), and therefore can be computed with a finite-dimensional Hermitian $\begin{bmatrix}H\end{bmatrix}$—even when the matrix is Hermitian. Equivalently, in the momentum space, the singular functions in solving those equations with $\mathrm{Im}\,E\to 0^{+}$ are smeared out when $\mathrm{Im}\,E$ is nonzero.
@@
After attaining converged calculations for the smallest $\mathrm{Im}\,E$ with given computing resources, our second step is to analytically continue the computed $\mathcal{A}(E)$ to the other domains of the complex $E$ plane, including the real axis. One main goal of this work is to develop RBM-based emulators to perform such analytical continuation. As demonstrated later, this approach effectively moves the BCTs below the real energy axis and could expose resonance states as eigenstates. It is closely connected to the so-called (near)-optimal-rational-approximation studied recently Trefethen et al. (2021); Trefethen (2023). The other goal of this work is to use RBM-based emulators to efficiently interpolate this analytical continuation in the $\bm{\theta}$ space.
@@
The strategy of utilizing those complex-$E$ calculations has already been employed elsewhere. Its simplest form would be a quick remedy used extensively in linear response theory during response function calculations Nakatsukasa et al. (2016). We use $-\mathrm{Im}\,\mathcal{A}(E_{0})$ with real $E_{0}$ as a response-like function to explain the procedure. As just discussed, a straightforward calculation with a Hermitian $H$ matrix produces $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ as a summation of $\delta$ functions on the real energy axis. The remedy is to smear out $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ by adding a finite imaginary part $i\eta$ to $E_{0}$—believed to be due to the so-called damping factor. This amounts to $-\mathrm{Im}\,\mathcal{A}(E_{0})\approx-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0}+i\eta)\end{bmatrix}$. From Fig. 1, it is reasonable to assume $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0}+i\eta)\end{bmatrix}\approx-\mathrm{Im}\,\mathcal{A}(E_{0}+i\eta)$. I.e., according to the remedy, it is better to assume $-\mathrm{Im}\,\mathcal{A}(E_{0}+i\eta)\approx-\mathrm{Im}\,\mathcal{A}(E_{0})$—meaning the exact result wouldn’t change from $E_{0}+i\eta$ to $E_{0}$, than using $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ directly. On the one hand, this shows the transition of the continuum physics from $E_{0}+i\eta$ to $E_{0}$ is smoother or simpler than what is produced by a large-matrix-based calculation. On the other hand, it points out the need for an extrapolation in $E$ that is at least better than assuming no change in that extrapolation.
@@
The other more sophisticated methods of this type include the CE method Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and the LIT methods Efros (1985a); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024). The two types of calculations extrapolate the complex-$E$ results to the real-$E$ axis using their specific procedures. Moreover, the $\bm{\theta}$ emulation discussed in this work could be helpful for the CE and LIT calculations. Sec. VI will go into detail about these aspects.
@@
II.3 Exponentially pole clustering near branch points and discretized BCTs in (near)-optimal rational approximation
@@
Lloyd N. Trefethen and others, in recent years, have done a series of numerical studies on the rational approximations of univariate functions, $f(z)$, with singularities, including branch points. A good review can be found in Ref. Trefethen (2023). In particular, they Trefethen et al. (2021) noticed that to achieve an efficient (or near-optimal) approximation, the poles of the rational approximation cluster exponentially near the branch points of $f(z)$, such as our $\mathcal{A}(z)$. Such pole distribution has been known in the case of approximating $|z|$ function (or equivalently $\sqrt{z}$ewman (1964) for some time, as mentioned in Ref. Trefethen et al. (2021). Still, these recent works emphasize the “widespread” of this pole clustering phenomenon in the near-optimal rational approximations. To have a visual impression of the clustering pattern of the poles or eigenenergies, see, for example, Figs. 3 and  4.
@@
These works also found that the distribution of the logarithms of the modulus between poles and the corresponding branch points is not precisely uniform. Instead, tapering off that distribution to zero when approaching the branch points is needed for more efficient approximation. Through numerical experimentation with various functions having branch points and heuristic arguments in the limit of dense distribution of the training (or interpolating) points, they show that the approximation errors generally scale as $e^{-c\sqrt{N_{\mathrm{eff}}}}$ with $N_{\mathrm{eff}}$ as the number of effective poles and $c$ being positive.
@@
Since the rational approximation performs the worst in the region where poles are located, those poles should lie away from the region of the training (or interpolating) data. Heuristically speaking, the area of the poles should be minimized to further reduce the errors in the complex plane. Therefore, when approximating a function with branch points, the poles “line up” to form discretized BCT lines, another general behavior observed and supported with some general understandings in these recent studies (see Chap. 7 in Ref. Trefethen (2023)). However, rigorously speaking, a rational function doesn’t have branch points and thus no BCTs. I.e., although the definition of the true BCTs is a convention for the exact function $f(z)$, the discretized BCTs emerge such that $f$’s rational approximation achieves the best performance. Effectively, we can consider (near)-optimal rational approximation chooses its own BCTs for the $f(z)$, a narrative we will use.
Lloyd N. Trefethen et al. have developed the AAA rational approximation algorithm Nakatsukasa et al. (2018, 2023) [^10] In this work, we utilize its Python implementation in the baryrat package Hofreither (2021).. It can be viewed as a “data-driven” emulator Duguet et al. (2024), as it only requires training data $f(z)$ at training (interpolating) points $z_{j}$ without knowing the underlying physics. It has been used in recent quantum physics studies Xu et al. (2022); Huang et al. (2023) [^11] Broadly speaking, the rational approximations have been used in quantum physics for analytically continuing the many-body Green’s function in the complex energies or frequencies to the real domain Goswami et al. (2024); Ying (2022)..
In the AAA approximation, $f(z)$ is approximated by the barycentric representation,
$r(z)=\frac{n(z)}{d(z)}=\left.\sum_{j=1}^{m}\frac{w_{j}f_{j}}{z-z_{j}}\middle/\sum_{j=1}^{m}\frac{w_{j}}{z-z_{j}}\right.\ .$ (7)
This representation of rational approximation provides numerical stability. Here $z_{j}$ are the training points and $f_{j}=f(z_{j})$. It is easy to see that $r(z_{j})=f_{j}$. To find the weights, $w_{j}$, the method minimizes the errors in terms of $f(z)d(z)-n(z)$, with other constraints. The method requires input about expected errors/noise in the data so that it doesn’t overfit and is stable against noise. As can be seen, $r(z)$ has only simple poles and approaches to constant at infinity. In contrast, Eq. (6) suggests that at $z\to\infty$, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$s in the NHQM approximations, including our emulator-based approach, goes to zero at the limit. However, this difference is not substantial, as we only focus on a finite domain, and poles far away in all these approximations effectively behave as polynomials of $z$. Also, note that the AAA algorithm provides the positions of its poles.
As will be shown through the numeric results in Secs. IV and V, there are intriguing similarities between our CEE results and the near-optimal-rational approximations concerning the location of the BCTs, the distribution of the discretized BCT poles, the approximation of $\mathcal{A}$, and the scaling of the approximation errors in terms of $N_{\mathrm{eff}}$. Specifically, we will compare our RBM-based emulators and the AAA approximations in several cases. It is tempting to equate the optimality of the rational approximation with the effectiveness of the RBM subspace projection in constructing the ROMs for a single variable.
It is worth pointing out that the eigenenergies of the NHQM calculations don’t have to line smoothly to form discretized BCTs in the complex $E$ plane, in addition to the bound and resonance states. Numerical noise and potentially lack of state density scatter these poles and distort the pattern. This does not conflict with the discussion of the general properties of the near-optimal rational approximations, as the NHQM calculations are not necessarily the optimal rational approximations of $\mathcal{A}$. However, the eigenenergies of the DC states must be well separated in the complex $E$ plane from the physical states, including resonances. Otherwise, it is impossible to identify the resonance states by only knowing their eigenenergies; as intimately connected, the $\mathcal{A}$ at real energies would have the DC states’ contributions comparable to those of the physical resonance poles, which gives us unphysical results and forces us back to the conundrum that the NHQM methods aim to help us avoid.
@@
In light of this discussion, we advocate using the pole separation as a basic test about whether the NHQM calculations capture necessary continuum physics: the resonance pole generates the correct resonance behavior for $\mathcal{A}$ at real energies; in contrast, the other poles (i.e., the discretized BCT) collectively produce the smooth background underneath the resonance peaks. For the latter, the near-threshold behavior inevitably requires enough density of the poles in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ towards the thresholds to achieve adequate approximations [^12] The definition of an adequate approximation will depend on how close to the threshold we want the calculation to be correct. A rational approximation of $\mathcal{A}$, to which any finite matrix calculations belong, has increasing errors towards the branch points because the approximant is infinitely differentiable there, but the underlying function is not. However, the problematic regions keep shrinking when the number of effective poles increases towards the branch points. This illustrates the challenge of a near-threshold calculation if we don’t build the known analytical structure into the calculation at its start., but the distributions of the DC state poles in other regions away from the branch points are not understood.

### v2

@@ II.2 NHQM methods
Figure 1: A resolvent matrix element $\mathcal{A}(E)$ in the complex $E$ plane with fixed sources. Here, the off-shell behavior of a $T$-matrix, specifically its non-Born term, is explored. The two rows show the real and imaginary parts of the amplitude. From left to right, three different calculations are shown. The first two are based on bound-state-like calculations, where I enforce a Dirichlet boundary condition on the relative motion, forcing the wave functions to be zero at $r=R_{IR}$; the third is the exact calculation, corresponding to the $R_{IR}\to\infty$ limit.
To illustrate the NHQM methods, including the one developed here, I use a two-body toy model with finite-range interactions. The toy system mimics a two-nucleon system in $s$-wave, for which I compute $\mathcal{A}$ in various ways. The sources are chosen such that $\mathcal{A}$ corresponds to the non-Born term in the scattering $T$-matrix, previously discussed as one option in Sec. II.1. My focus here is on the generic features of the numerical results, while the numeric details can be found in the beginning of Sec. IV.
@@
The figure plots the real (top row) and imaginary (bottom row) parts of various $\mathcal{A}(E)$s in the complex $E$ plane; from left to right, the first two columns are from the $\mathcal{A}$ calculations where finite $H$ matrices are inserted in Eq. (1). To get such $H$ matrices, I enforce a boundary condition such that all the states are zero at $r=R_{IR}$ (10 and 50 fm in the 1st and 2nd column), a particular long-distance (infrared, IR) regulator, while a series of short-distance (ultra-violet, UV) regulators  [^5] In a finite basis, such as a coordinate-space Lagrange mesh [80] used in my two-body calculations or a plane-wave basis, the resolution of the small-distance scale, i.e., mesh-point spacing or momentum cut-off in the respective bases, is finite, which is known as a short-distance (UV) regularization. In general, this resolution improves with basis size. are tested so that the results shown in the figure converge with respect to increasing the UV resolution. The last column shows the exact results, essentially $\mathcal{A}$’s limit at $R_{IR}\to\infty$.
The sharp features around the real axis, as seen in the first two columns, can be understood via the spectrum of the $H$ matrices. Specifically, the approximated $\mathcal{A}$ with finite $R_{IR}$ and UV cut-off is a finite sum of poles if one consider it as a function of $E$:
@@
$|\Psi_{i}\rangle$ is the $i$th eigenstate of $\begin{bmatrix}H\end{bmatrix}$ and the associated eigenenergy $E_{i}$ determines a pole location. Note here and later, I label the finite-matrix-based approximations of a quantity, e.g., $\mathcal{A}$, by $\begin{bmatrix}\mathcal{A}\end{bmatrix}$. Since $\begin{bmatrix}H\end{bmatrix}$ is kept Hermitian while I construct a basis to project $H$, the poles are only on the real axis.
@@
One would have three observations: (1) above and below the real axis, the exact $\mathcal{A}$ (in the 3rd column) is smooth almost every where, except the $E=0$ point where the function is not as well behaved. When crossing the positive real axis, the amplitude experiences a discontinuity. This is a reflection of $E=0$ being $\mathcal{A}$’s branch point and the positive real axis as the BCT; (2) when taking $R_{IR}\to\infty$ limit, the poles to the right of the $E=0$ point become more and more densely distributed and eventually merge into a BCT along the real axis; and (3) to the left of $E=0$, there are no densely distributed poles except possible isolated bound-state poles (absent in this toy model). $E=0$ is such a nontrivial location because it is the model’s continuum threshold, below which the kinetic phase space for the scattering eigenstates is zero and above which the phase space starts increasing from zero [^6] This can also be visualized by plotting $\mathrm{Im}\,\mathcal{A}(E)$ with $\mathrm{Im}\,E\to 0^{+}$. The function is zero below the threshold and grows starting at the threshold. For example, see the 3rd panel in Fig. 12..
That is, at $R_{IR}\to\infty$ limit, the summing over the index $i$ in Eq. (5) turns into integrating over the continuous eigenenergy variable $\tilde{E}$:
$\mathcal{A}$ $=\sum_{B}\frac{\langle\tilde{S}|\Psi_{B}\rangle\langle\Psi_{B}|S\rangle}{E-E_{B}}$
$+\int_{-\infty}^{\infty}d\tilde{E}\frac{\langle\tilde{S}|\Psi_{\tilde{E}}\rangle\langle\Psi_{\tilde{E}}|S\rangle}{E-\tilde{E}}\Theta(\tilde{E}-E_{\text{th}})\ .$ (6)
Here, $\Psi_{B}$ and $\Psi_{\tilde{E}}$ are, respectively, the bound state in the discrete spectrum and the other eigenstates in the continuous spectrum; $E_{\text{th}}$ is the location of the threshold (located at 0 here). It is the integration support defined by the step function $\Theta(\tilde{E}-E_{\text{th}})$ that gives rise to the branch point at $E=E_{\text{th}}$. Again, the BCT defined along the real axis is due to the continuous real eigenenergies.
@@
The above discussions can be generalized to more complex situations where multiple thresholds (or channels) exist. At infinite-$R_{IR}$ limit, there are the same number of eigenstate groups as the number of thresholds [^7] The eigenstates of the full Hamiltonian provide a complete basis to expand the full space. See Chap. 7 in Ref. [77]. At each threshold, a new group of eigenstates emerges as initiated by the corresponding incoming channel.. Each group starts to contribute to $\mathcal{A}$ above the corresponding threshold (labeled by the associated channel $c$) in a simple additive way, i.e., the integration in Eq. (6) can be broken into individual groups, each with an integral support $\Theta(\tilde{E}-E_{\text{th},c})$. Repeating the discussion in the single-channel case, one see that $\mathcal{A}(E)$ has a branching point at each $E=E_{\text{th},c}$. It is worth pointing out that at finite $R_{IR}$, the discussion could be tricky, as each discrete eigenstate can’t simply be labeled by a particular incoming channel $c$.
In addition, it will be interesting and also possible to generalize these conclusions to the cases with long-range repulsive Coulomb interactions (needed in nuclear physics), in which the branch points are also essential singularities due to the Coulomb-barrier-induced Gamow suppression factors.
@@
To understand the failure of the finite matrix calculations, one returns to the first two columns in Fig. 1. The regions with the largest errors are around the real energy axis, where the poles are located [^8] It is worth pointing out that the discretized BCT poles can be partially understood through quantization conditions for the corresponding eigenvalues. At least for simple systems, the continuum physics can be extracted at the poles’ locations using Lüscher-type approaches [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95]. These methods also extract the continuum physics by computing bound-state-like problems.. That is, unfortunately, where the continuum physics observables can be physically measured. At the same time, the calculations potentially already converge in the other regions. This is the essential feature of the rational approximation [16], to which Eq. (5) belongs, applied to a function with branch points: poles are assigned to mimic branch cut so that the approximation works in the region away from the BCTs but not around the BCTs. Therefore, the cause of the continuum-physics difficulty is the Hermiticity of the $\begin{bmatrix}H\end{bmatrix}$, although Hermitian Hamiltonians have been used extensively in computing discrete bound states.
@@
Existing NHQM methods, such as the complex scaling method [10, 11, 14, 12, 13], solve this problem by giving up $\begin{bmatrix}H\end{bmatrix}$’s Hermiticity so that the BCTs and the problematic region of the rational approximation are moved away from the real axis. As a result, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ can converge with finite $R_{IR}$ around the real energy region. Moreover, if the BCT is moved further away enough that the resonance poles, which were on the second Riemann sheet [^9] A second Riemann sheet is sometimes called as an unphysical sheet, which could lead to the wrong impression that it is not important. On the contrary, the function’s behavior in that part of the complex plane is directly related to the function’s behaviors relevant to observations. when the BCTs are defined as along the real axis, are now in the region enclosed by the old and new BCTs. I.e., these resonance poles become the poles on the 1st Riemann sheet. Therefore, the resonance poles can now be seen simultaneously in $\mathcal{A}$ with bound state poles. Eq. (1) then dictates that the resonances, as the bound state, are part of the $H$’s eigenstates. In addition, all the eigenstates, including the bound and resonance states and the so-called discretized continuum (DC) states that give rise to the discretized BCT poles in $\mathcal{A}$, are integrable. Therefore, the NHQM methods enable computing the continuum physics from the bound-state-like calculations.
@@
The other solution, which I follow here, is based on the observation that $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$ can converge to a certain required precision when the dimension of the $\begin{bmatrix}H\end{bmatrix}$ and the size of $|\mathrm{Im}\,(E)|$ are large enough. As shown in Fig. 1, with $R_{IR}=10$ fm, the computed $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ with $|\mathrm{Im}\,E|>10$ MeV already converges nicely to the exact results. This statement applies to general cases that the wave function solutions in Eqs. (2) and (3) for complex $E$ and compact sources satisfy the bound-state-like boundary conditions [18], and therefore can be computed with a finite-dimensional Hermitian $\begin{bmatrix}H\end{bmatrix}$—even when the matrix is Hermitian. Equivalently, in the momentum space, the singular functions in solving those equations with $\mathrm{Im}\,E\to 0^{+}$ are smeared out when $\mathrm{Im}\,E$ is nonzero.
@@
After attaining converged calculations for the smallest $\mathrm{Im}\,E$ with given computing resources, my second step is to analytically continue the computed $\mathcal{A}(E)$ to the other domains of the complex $E$ plane, including the real axis. One main goal of this work is to develop RBM-based emulators to perform such analytical continuation. As demonstrated later, this approach effectively moves the BCTs below the real energy axis and could expose resonance states as eigenstates. It is closely connected to the so-called (near-)optimal-rational-approximation studied recently [96, 16]. The other goal of this work is to use RBM-based emulators to efficiently interpolate this analytical continuation in the $\bm{\theta}$ space.
@@
The strategy of utilizing those complex-$E$ calculations has already been employed elsewhere. Its simplest form would be a quick remedy used extensively in linear response function calculations (see e.g., [97]). Here, I use $\mathrm{Im}\,\mathcal{A}(E_{0})$ with real $E_{0}$ as a response-like  [^10] The term “response-like” emphasized that $\mathrm{Im}\,\mathcal{A}(E_{0})$ is generally not a physical response function (unless the sources are properly chosen), but they do share the same structure as the imaginary part of a resolvent matrix element evaluated at $E_{0}+i0^{+}$. function to explain the procedure. As just discussed, a straightforward calculation with a Hermitian $H$ matrix produces $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ as a summation of $\delta$ functions on the real energy axis. The remedy is to smear out $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ by adding a finite imaginary part $i\eta$ to $E_{0}$—believed to be due to the so-called damping factor. This amounts to $\mathrm{Im}\,\mathcal{A}(E_{0})\approx-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0}+i\eta)\end{bmatrix}$. From Fig. 1, it is reasonable to assume $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0}+i\eta)\end{bmatrix}\approx-\mathrm{Im}\,\mathcal{A}(E_{0}+i\eta)$. Therefore, the remedy essentially assumes that $\mathrm{Im}\,\mathcal{A}(E_{0}+i\eta)\approx-\mathrm{Im}\,\mathcal{A}(E_{0})$, i.e., the exact result $\mathrm{Im}\,\mathcal{A}(E)$ wouldn’t change when it is extrapolated from $E=E_{0}+i\eta$ to $E=E_{0}$. Empirically, such extrapolation works better than using $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}(E_{0})\end{bmatrix}$ directly! On the one hand, this shows the transition of the continuum physics from $E_{0}+i\eta$ to $E_{0}$ is smoother or simpler than what is produced by a large-matrix-based calculation. On the other hand, it points out the need for an extrapolation in $E$ that is at least better than assuming no change in that extrapolation.
@@
The other more sophisticated methods of this type include the CE method [17, 58, 59, 60, 61, 62, 63, 64, 65] and the LIT methods [18, 66, 67, 68, 69, 70, 71]. The two types of calculations extrapolate the complex-$E$ results to the real-$E$ axis using their specific procedures. Moreover, the $\bm{\theta}$-emulation discussed in this work could be helpful for the CE and LIT calculations. Sec. VI will go into detail about these aspects.
@@
II.3 Exponentially pole clustering near branch points and discretized BCTs in (near-)optimal rational approximation
Lloyd N. Trefethen and others, in recent years, have done a series of numerical studies on the rational approximations of a univariate function ($f(z)$) with branch points and other singularities, which approximates $f$ as a sum of simple poles:
$f(z)\approx r(z)=\sum_{i=1}^{N_{p}}\frac{w_{i}}{z-z_{i}}\ ,$ (7)
with $z_{i=1,2,,,N_{p}}$ as the locations of the poles. One of the poles could be located at $\infty$—meaning a polynomial could be added here.
@@
A good review can be found in Ref. [16]. In particular, they [96] noticed that to achieve an efficient (or near-optimal) approximation, the poles cluster exponentially near the branch points of $f(z)$, such as the $\mathcal{A}(z)$. Such pole distribution has been known in the case of approximating $|z|$ function (or equivalently $\sqrt{z}$) [98] for some time, as mentioned in Ref. [96]. Still, these recent works emphasize the “widespread” of this pole clustering phenomenon in the near-optimal rational approximations. To have a visual impression of the clustering pattern of the poles or eigenenergies, see, for example, Figs. 3 and 4.
@@
These works also found that the distribution of the logarithms of the modulus between poles and the corresponding branch points is not precisely uniform. Instead, tapering off that distribution to zero when approaching the branch points is needed for more efficient approximation. Through numerical experimentation with various functions having branch points and heuristic arguments in the limit of dense distribution of the training (or interpolating) points, they [96, 16] show that the approximation errors generally scale as $e^{-c\sqrt{N_{p}}}$ with $c$ being positive.
@@
Since the rational approximation performs the worst in the region where poles are located, those poles should lie away from the region of the training (or interpolating) data. Heuristically speaking, the area of the poles should be minimized to further reduce the errors in the complex plane. Therefore, when approximating a function with branch points, the poles “line up” to form discretized BCT lines, another general behavior observed and supported with some general understandings in these recent studies (see Chap. 7 in Ref. [16]). However, rigorously speaking, a rational function doesn’t have branch points and thus no BCTs. I.e., although the definition of the true BCTs is a convention for the exact function $f(z)$, the discretized BCTs emerge such that $f$’s rational approximation achieves the best performance. Effectively, one can consider (near-)optimal rational approximation chooses its own BCTs for the $f(z)$, a narrative I will use.
@@
Note that the definition of an adequate approximation depends on how close to the threshold one wants the calculation to be correct. A rational approximation of $\mathcal{A}$, to which any finite-matrix-based calculations belong, has increasing errors towards the branch points because the approximant is infinitely differentiable there, but the exact function is not. However, the problematic regions keep shrinking when the number of effective poles increases towards the branch points. This illustrates the challenge of a near-threshold calculation if one don’t build the known analytical structure into the calculation at its start.
A so-called AAA algorithm has been used in these demonstrations [99, 100] [^11] In this work, I utilize its Python implementation in the baryrat package [101]., which is considered as (near-)optimal rational approximation [96]. It can be viewed as a “data-driven” emulator [2], as it only requires training data $f(z)$ at training (interpolating) points $z_{j}$ without knowing the underlying physics. It has been used in recent quantum physics studies [102, 103], for example, to analytically continue the many-body Green’s function in the complex energies or frequencies to the real domain [104, 105].
In this algorithm, $f(z)$ is constructed in the barycentric representation of Eq. (7):
$r(z)=\frac{n(z)}{d(z)}=\left.\sum_{j=1}^{m}\frac{w_{j}f_{j}}{z-z_{j}}\middle/\sum_{j=1}^{m}\frac{w_{j}}{z-z_{j}}\right.\$. (8)
This representation provides numerical stability. Here $z_{j}$ are the training points and $f_{j}=f(z_{j})$. It is easy to see that $r(z_{j})=f_{j}$. To find the weights, $w_{j}$, the method minimizes the errors in terms of f(z)d(z)n(z)f(z)d(z)-n(z), with other constraints. The method requires input about expected errors/noise in the data so that it doesn’t overfit and is stable against noise. As can be seen, $r(z)$ has only simple poles and approaches to constant at infinity. In contrast, Eq. (5) suggests that at $z\to\infty$, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$s in the NHQM approximations, including my emulator-based approach, goes to zero at the limit. However, this difference is not substantial, as I only focus on a finite domain, and poles far away in all these approximations effectively behave as polynomials of $z$. Also, note that the AAA algorithm provides the positions of its poles.
As will be shown through the numeric results in Secs. IV and V, there are intriguing similarities between my CEE results and the (near-)optimal-rational approximations concerning the location of the BCTs, the distribution of the discretized BCT poles, the approximation of $\mathcal{A}$, and the scaling of the approximation errors in terms of the number of poles. Specifically, I will compare my RBM-based emulators and the AAA approximations in several cases. It is tempting to equate the optimality of the rational approximation with the effectiveness of the RBM subspace projection in constructing the ROMs for a single variable.